# Quasi-randomization tests for network interference

### v1

@@ 7   Conclusion
In this paper, we develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup. Our proposed method assumes mild distributional conditions in the network-generating process to build a quasi-randomization test. Whilst promising, there are several open problems that remain. Incorporating covariates or nodal attribute information to better characterize the random graph null model may help enhance the power of the test and is an interesting direction for future work. We also assume fully and correctly specified information of the given network. An analysis of the random graph null models under partial information and network misspecification is another interesting avenue for future work. Finally, our focus in this paper has been on randomized experiments exclusively. Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and important extension is to study random graph null testing procedures in observational studies.
Acknowledgments
We thank Sumit Kunnumkal, Parshuram Hotkar, and ISB seminar participants for their helpful comments and feedback. We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access. Pallavi Basu’s research is partially supported by the SERB MATRICS award MTR/2022/000073. There is no conflicting interest to declare.
@@ Appendix A Algorithm: $\epsilon$-net cluster
  • Step 1: Pick the first vertex in $S$, say $v_{1}$
  • Step 2: Find all units within $\epsilon$ distance of $v_{1}$
@@
Output: C
@@ B.1   Proof of Theorem 1:
$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}),$
Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z_{obs})$. Thus,
@@
$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
@@ B.2   Proof of Theorem 2:
$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
@@
We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a cluster randomized experiment. Consider $g,g^{\prime}\in[G_{obs}]$.
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}])\cdot P(G=g|G\in[G_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}])},$ (9)
@@
$P(Z=Z_{obs}|G=g,G\in[G_{obs}])$ $=P(\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\hskip 5.69054ptj\in[N]|G=g,G\in[G_{obs}]),$ (11)
$=P(W=W_{obs}|G=g,G\in[G_{obs}]),$
$=P(W=W_{obs})$
$=P(W=W_{obs})|G=g^{\prime},G\in[G_{obs}]),$
@@
Note that we take cluster partition to be fixed as part of the experimental design. Using Equation (10) and (11) in Equation (9), we obtain
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}]),$
Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under cluster randomized experiment. Hence,
@@ Appendix C Additional numerical simulations
We present the data-generating process considered in the simulations. To give an overview, we consider a linearly separable potential outcome model for the outcome function, which is a composition of two effects: a direct treatment effect and a spillover treatment effect. The spillover parameter is weighted by a function of the unit’s neighbor. We consider two such outcome models for the simulations: one that captures spillovers by the presence of a treated neighbor and the other that captures it via the proportion of treated neighbors.
C.1   Data generating mechanism
@@
$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}>0)+\epsilon_{i},$ (12)
@@
The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and giving them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). Here, $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have 100 units and 5 treated units in the model in one case. We then extended the study to 500 units, and 250 were treated in the second case.
@@
    Direct treatment effect: We take $\tau_{direct}=4$ in our simulations. For robustness, we also perform the simulations with $\tau_{direct}=1$. It is noted that $\tau_{direct}$ has no impact on the validity and power of the test.
@@
    Spillover effect: We vary the spillover effect, $\tau_{spill}$ from 0 to 7, keeping $\tau_{direct}$ fixed.
@@
    Test statistic: We consider the test statistic taking the difference of treatment effect on control units with control neighbors and control units with treated neighbors, $T_{I_{c}}(G|Z_{obs})$. We also consider the test statistic based on the difference in treatment effect of a low proportion of treated neighbors and a high proportion of treated neighbors, $T_{{quant}_{c}}(G|Z_{obs})$. Refer to Section 4.3 for further details.
We generate the randomization distribution of the test statistic in consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2. The simulations in Table 3 were carried out on 100 units, with 5 being treated and the rest in control. We approximated the randomization distribution of the test statistic by taking 200 samples of the graph. We calculate the corresponding p-value by the proportion of times the test statistic imputed was larger than the test statistic for the observed network. If the p-value is less than 0.05, we reject the null hypothesis of no spillover effects. We replicated the above 480 times and calculated rejection rates as the proportion of times the p-value was significant. For this setup, the outcome model (12) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 3.
We then repeat the simulations for the above setup for outcome model (13) and present them in Table 4.
We note that the power of the test decreases substantially for model (13), compared with model (12). This is owing to the simplistic nature of the outcome model (12), and we present its results for verification purposes. Another caveat in the above setting is the nature of test statistic $T_{I_{c}}(G|Z_{obs})$. Depending on the input graph, we might not get any control units with all neighboring ones in control, invalidating the procedure. So, the above setup involves 100 units with 5 units in treatment. We numerically found the procedure failing when the number of treated units was increased to more than 5. We present the results for model (13) with $T_{{quant}_{c}}$ as test statistic in Table 4. Here, the number of units in the population is 500, with the number of treated units being 250. $\tau_{direct}$ is 4 and varying $\tau_{spill}$ is used for power performance measure.
In all the above models, we took the probability of an edge between two nodes as $0.2$. We present the same setup in 4, with the probability of an edge between two nodes estimated as below. $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.
@@
Here, $|E_{G_{obs}}|$ represents number of edges in the observed graph $G_{obs}$. The results are presented in Table 4.
We highlight how type 1 error is controlled under 0.05 in all the setups (case when $\tau_{spill}$ is zero). When the null hypothesis is not true, we observe good power against the test in line with results from Athey et al. [2018]. This remains true when the network distribution is estimated parametrically.
Table 3: Rejection rates for null hypothesis of no spillover effects; Outcome model (12) with test statistic $T_{I_{c}}$
@@
0.0 0.035
0.1 0.083
0.2 0.173
0.3 0.279
0.4 0.402
0.5 0.579
0.6 0.745
0.7 0.887
0.8 0.943
0.9 0.987
1.0 0.993
Table 4: Rejection rates for null hypothesis of no spillover effects; Outcome model (13) with test statistic $T_{I_{c}}$, $T_{{quant}_{c}}$ and $T_{{quant}_{c}}$ with estimated Erdős-Rényi random graph
@@
$T_{I_{c}}$ $T_{{quant}_{c}}$ $T_{{quant}_{c}}$ with $\hat{p}$
0 0.054 0.040 0.054
1 0.058 0.080 0.069
2 0.110 0.238 0.229
3 0.181 0.419 0.419
4 0.256 0.694 0.693
5 0.365 0.881 0.881
6 0.485 0.954 0.954
7 0.612 0.990 0.994

### v2

@@ 7 Conclusion
We develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup. Our proposed method assumes natural distributional conditions in the network-generating process to build a quasi-randomization test. While promising, several open problems remain. Incorporating covariates or nodal attribute information to characterize the random graph null model better may help enhance the power of the test and is an interesting direction for future work. We also assume fully and correctly specified information about the given network. Analyzing the random graph null models under partial information and network misspecification is another exciting avenue for future work. Finally, our focus in this paper has been on randomized experiments exclusively. Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and essential extension is to study random graph null testing procedures in observational studies.
Acknowledgments. We thank Avi Feller, Fabrizia Mealli, Panos Toulis, ACIC 2024 conference participants, and the Tom Ten Have Award committee for their encouragement and feedback. We also thank Sai Sriramya Gorripati, Parshuram Hotkar, Sumit Kunnumkal, and ISB seminar participants for their enthusiasm and comments. We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access.
@@ Appendix A Algorithm: $\epsilon$-net cluster
  • Step 1: Pick the last vertex in $S$, denoted by $v$, and store in $V$
  • Step 2: Find all units within $\epsilon$ distance of $v$
@@
Permute the outputs to the original ordering and store them in $C,V$
Output: C and V
@@ B.1 Proof of Theorem 1:
$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}).$
Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution induced by $P(G|Z_{obs})$. Thus,
@@
$T_{c}(G|Z_{obs})\stackrel{{\scriptstyle d}}{{=}}T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0}.$
@@ B.2 Proof of Theorem 2 and Theorem 3:
$T_{c}(G|Z_{obs})\stackrel{{\scriptstyle d}}{{=}}T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0}.$
@@
Let $[G_{obs}]=\{G\in\{0,1\}^{N\times N}:\>deg(G)=deg(G_{obs})\}$. We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a completely randomized experiment. Consider $g,g^{\prime}\in[G_{obs}]$.
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}])\cdot P(G=g|G\in[G_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}])}.$ (8)
@@
$P(Z=Z_{obs}|G=g,G\in[G_{obs}])$ $=\frac{1}{\binom{N}{\sum Z_{obs}}},$ (10)
@@
Using Equation (9) and (10) in Equation (8), we obtain
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}]).$
Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under completely randomized experiment. Therefore,
@@
Hence, we prove that Procedure $2$ is valid. We now prove the validity of Procedure $3$. Let
$[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\>|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
Here, ${Z_{obs}}^{t}=\{i\in\mathbb{P}:\>{Z_{obs}}_{i}=t\}$ for $t\in\{0,1\}$. Note that $[G_{obs}|Z_{obs}]\subseteq[G_{obs}]$, where $[G_{obs}]=\{G\in\mathcal{G}|G\cong G_{obs}\}$. We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])$ is a uniform distribution under a cluster randomized experiment. Let $C$ be a random variable defined as $\mathcal{A}\circ G$. Here, $\mathcal{A}$ represents the clustering algorithm in Algorithm 1.
Consider $g,g^{\prime}\in[G_{obs}|Z_{obs}]$. Then $P(G=g|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])=$
$\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}|Z_{obs}])\cdot P(G=g|G\in[G_{obs}|Z_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}|Z_{obs}])}.$ (11)
$P(G=g|G\in[G_{obs}|Z_{obs}])$ $=\frac{f_{\eta}(s(g))}{\sum_{G\in[G_{obs}|Z_{obs}]}f_{\eta}(s(G)},$ (12)
$=\frac{f_{\eta}(s(\pi(G_{obs})))}{\sum_{G\in[G_{obs}|Z_{obs}]}f_{\eta}(s(\pi^{\prime}(G_{obs}))},$
$=\frac{f_{\eta}(s(G_{obs}))}{\sum_{G\in[G_{obs}|Z_{obs}]}f_{\eta}(s(G_{obs})},$
$=P(G=g^{\prime}|G\in[G_{obs}]).$
$P(Z=Z_{obs}\>|\>G=g,$ $G\in[G_{obs}|Z_{obs}])$ (13)
$=P(\sum_{i=1}^{|C|}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\;j\in[N]\>|\>G=g,G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|\pi(C_{obs})|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))={Z_{obs}}_{j}\;j\in[N]\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(\pi^{-1}(j)\in{C_{obs}}_{i})=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in{C_{obs}}_{i})=t\quad\forall\pi(j)\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
$=P(W=W_{obs})=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}|Z_{obs}]).$
Using Equation (12) and (13) in Equation (11), we obtain
$P(G=g|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}|Z_{obs}]).$
Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}|Z_{obs}]\}$ is uniform under cluster randomized experiment. Therefore,
$p_{Z_{obs},G_{obs}}=\frac{1}{|[G_{obs}|Z_{obs}]|}\cdot\sum_{G\in[G_{obs}|Z_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
Hence, we prove that Procedure $3$ is valid. ∎
@@ Appendix C Additional numerical simulations
We present the Erdős-Rényi random graph model and the data-generating process considered in the simulations here. To give an overview, we consider a linearly separable potential outcome model for the outcome function, composed of two effects: a direct treatment effect and a spillover treatment effect. The spillover parameter is weighted by a function of the unit’s neighbor. We consider two such outcome models for the simulations: one that captures spillovers by the presence of a treated neighbor and the other that captures it via the proportion of treated neighbors.
C.1 Erdős-Rényi random graph model
We formally describe the Erdős-Rényi random graph model (Erdős et al. [1960]). The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model. We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$. That is,
To model $P(G)$ using Equation (14), we have to estimate the parameter $p$. We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$. We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test. We also obtain the power of the test under different scenarios. An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world. For example, one observes a high degree of clustering in social networks in the graph (Watts and Strogatz [1998]). One drawback of the model is the independence and identical probability of formation of two edges in the graph. One way to characterize this is to evaluate the degree distribution of the random graph. It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution. Most real-life networks violate this property and have highly skewed degree distribution. In the main text, we generalize this and characterize random graphs with arbitrary degree distribution.
C.2 Data generating mechanism
@@
$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot z_{j}>0)+\epsilon_{i},$ (15)
@@
The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and giving them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). Here, $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have 100 total units and 5 treated units in the model in one case. We then extended the study to 500 units, and 250 were treated in the second case.
@@
    Direct treatment effect: We take $\tau_{direct}=4$ in our simulations. We also perform the simulations with $\tau_{direct}=1$ for robustness. It is noted that $\tau_{direct}$ has no impact on the validity and power of the test.
@@
    Spillover effect: We vary the spillover effect, $\tau_{spill}$ from 0 to 5, keeping $\tau_{direct}$ fixed.
@@
    Test statistic: We consider the test statistic taking the difference of treatment effect on control units with control neighbors and control units with treated neighbors, $T_{I_{c}}(G|Z_{obs})$. We also consider the test statistic based on the difference in treatment effect of a low proportion of treated neighbors and a high proportion of treated neighbors, $T_{quant}(G|Z_{obs})$. Refer to Section 4.3 for further details.
We generate the randomization distribution of the test statistic in consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2. The simulations in Table 3 were carried out on 100 units, with 5 being treated and the rest in control. We approximated the randomization distribution of the test statistic by taking 200 samples of the graph. We calculate the corresponding p-value by the proportion of times the test statistic imputed was larger than the test statistic for the observed network. If the p-value is less than 0.05, we reject the null hypothesis of no spillover effects. We replicated the above 450 times and calculated rejection rates as the proportion of times the p-value was significant. For this setup, the outcome model (15) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 3, varying from $0.0$ to $0.7$.
We then repeat the simulations for the above setup for outcome model (16) and present them in Table 4. We note that the power of the test decreases substantially for model (16), compared with model (15). This is owing to the simplistic nature of the outcome model (15), and we present its results for verification purposes. Another caveat in the above setting is the nature of test statistic $T_{I_{c}}(G|Z_{obs})$. We might not get any control units with all neighboring ones in control depending on the input graph, invalidating the procedure. So, the above setup involves 100 units with 5 units in treatment. We numerically found the procedure failing when the number of treated units was increased to more than 5. We present the results for model (16) with $T_{quant}$ as test statistic in Table 4. Here, the number of units in the population is 500, with the number of treated units being 250. $\tau_{direct}$ is 4 and varying $\tau_{spill}$ is used for power performance measure.
In all the above models, we took the probability of an edge between two nodes as $0.2$. We present the same setup in 4, with the probability of an edge between two nodes estimated below. $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.
@@
Here, $|E_{G_{obs}}|$ represents number of edges in the observed graph $G_{obs}$. The results are presented in Table 4. We highlight how type 1 error is controlled under 0.05 in all the setups (when $\tau_{spill}$ is zero). When the null hypothesis is false, we observe moderate power. This holds when the network distribution is estimated parametrically.
Table 3: Rejection rates for the null hypothesis of no spillover effects. Outcome model (15) with test statistic $T_{I_{c}}$.
@@
0.0 0.038
0.1 0.077
0.2 0.167
0.3 0.293
0.4 0.444
0.5 0.620
0.6 0.749
0.7 0.858
Table 4: Rejection rates for the null hypothesis of no spillover effects. Outcome model (16) with test statistic $T_{quant}$, and $T_{quant}$ with estimated Erdős-Rényi random graph.
@@
$T_{quant}$ $T_{quant}$ with $\hat{p}$
0 0.057 0.044
1 0.124 0.138
2 0.402 0.469
3 0.747 0.729
4 0.947 0.927
5 0.984 0.989
Appendix D Performance of test statistic $T_{bond}$ on Procedure 2
Figure 2: Distribution of p-values for a completely randomized experiment under Procedure 2 (isomorphism permutation variant) over 2,000 replications; the plots on the top have direct effect 0 with test statistics $T_{bond}$ on the left panel and $T_{quant}$ on the right. The plots on the bottom have a direct effect 4. All plots have a spillover effect of 0.4. The (blue) line indicates the threshold of 0.05. [Lower panel, left plot] The closer-to-uniform distribution of the p-values for $T_{bond}$ contributes to the drop in power as observed in Table 1.